A loop of wire in the shape of a rectangle rotates with a frequency of 219 rotation per minute in an applied magnetic field of magnitude 6 T. Assume the magnetic field is uniform. The area of the loop is A = 4 cm2 and the total resistance in the circuit is 7 Ω.

Find the maximum induced emf.
e m fmax =

Find the maximum current through the bulb.
Imax
Answer #1

angular velocity of loop, w = 219 rev/min

= 219*2*pi/60

= 22.93 rad/s

Area of the loop, A = 4 cm^2 = 4*10^-4 m^2

we know, maximum induced emf = N*A*B*w (here N is no of loops and B is magnetic field)

= 1*4*10^-4*6*22.93

= 0.055 volts or 55 mV <<<<<<---------Answer

maximum current through the bulb, Imax = emf_max/R

= 0.055/7

= 7.86*10^-3 A or 7.86 mA <<<<<<---------Answer